# The Preliminary Mauve Science Programme: Science themes identified for the first year of operations–ACRONYMS The Preliminary Mauve Science Programme: Science themes identified for the first year of operations

### v2

@@ Abstract
Pioneering a new model for science satellites and space science (Archer et al., 2020), Mauve launched on 28th November 2025 and will begin a collaborative scientific programme in early 2026, that focuses on monitoring stellar activity and variability. A major strength of the Mauve science programme is the availability of thousands of observational hours each year for its consortium members. This makes it ideal for time-domain astronomy, where hundred of hours are dedicated to long-duration observations and repeat observations for hundreds of targets. Additionally, Mauve will be ideally placed to conduct pilot studies and high-risk investigations.
@@ 2 Platform and Payload
As final operational performance will only be known post-commissioning, the BSSL team and the Mauve Science Team are considering multiple performance scenarios to prepare the science programme and optimise the observation schedule. MauveSim (Saba et al., 2025), an end-to-end simulator, was developed to assess the science capabilities of the satellite. The simulations presented here are derived from the most recent ground-based testing and expected in-orbit performance, with appropriate margins applied. The detector temperature, for instance, directly influences the thermally induced noise on the detector and the system’s overall performance. Following commissioning, MauveSim will be revised and updated to reflect the final in-flight performance. Pending commissioning and final performance, some Mauve science themes may be updated to account for spacecraft in-orbit behaviour.
@@ 3 Commissioning and Calibration
Mauve is equipped with a low-resolution spectrometer, resulting in each detector resolution element spanning approximately 40 pixels. This configuration leads to significant oversampling of the instrument line function – a matrix that describes photon energy redistribution, i.e., the probability of detecting an incident photon at any given pixel. Due to this oversampling, each pixel collects photons from a range of wavelengths, making it impossible to determine the exact energy of a photon striking a specific pixel. Any observation obtained with Mauve can be represented as the matrix product of the true stellar spectrum, the telescope’s effective area (ARF, representing total detector throughput), and the instrument line function (RMF). However, this matrix operation is non-reversible, meaning that the true stellar spectrum cannot be directly separated from the instrumental effects to recover the intrinsic stellar flux in physical units. Because all flux calibration information is contained within the ARF and RMF, calibrating Mauve involves adjusting these arrays such that the observed spectrum of a stable reference star with a well-known spectrum matches the simulated spectrum of the same source. When the observed and simulated spectra of the reference star match with each other, the calibration is deemed successful. If discrepancies remain, the ARF and RMF are iteratively refined until a satisfactory match is achieved. During the commissioning phase, and routinely throughout nominal operations, BSSL will generate and update the calibration files required by users to perform accurate data analysis.
@@
A comprehensive commissioning and calibration report will be released following Mauve’s commissioning period, summarising the lessons learned during spacecraft commissioning and outlining the achievable calibration performance of the instrument.
@@ 4.1 Mauve Science Planning
While Blue Skies Space Ltd. is responsible for developing the satellite platform and facilitating the planning process, the Mauve Survey Programme is a collaborative, member-led initiative. Survey members define the science themes, objectives, target lists, and observational strategies in coordination with Blue Skies Space Ltd., ensuring that scientific return is maximised through coordinated planning and transparent decision-making. The survey programme is led by the Mauve Science Team, a consortium of scientists from ten institutions across Europe, Asia, and North America, including:
@@ 4.2 Science Overview
The selected science themes can be broadly categorised into the following research areas: stellar activity and variability, host-star/planet interaction, hot stars, and exotic populations in binaries. Specific themes can then also be grouped by three main observation strategies, i.e. long-duration monitoring programmes, short-cadence single snapshots, and short-cadence monitoring with repeats. These categories determine scheduling strategy and observation cadence. Long-duration monitoring is typically focused on tracking stellar variability on specific targets over time. The themes using this observation strategy include studies of stellar activity and variability, specifically to understand the flare occurrence rates and to constrain the signatures of coronal mass ejections through flare-associated dimming.
The second category comprises short-cadence single snapshots, typically associated with large-scale statistical programmes observing many targets, or small groups of targets observed repeatedly. The themes related to short-cadence single-snapshot observations include investigations of exoplanet host stars and future Habitable Worlds Observatory (HWO) targets, the study of quiescent near-ultraviolet emission in low-mass stars to constrain the evolution of rotation and magnetic activity in Sun-like (and lower-mass) stars, and the classical Be star survey, designed to explore the mechanism that causes disk ejection.
@@ 4.3.1 Radiation mechanisms of stellar flares on M-dwarfs and young Sun-like stars
The UV and X-ray radiation from stellar flares is thought to influence planetary atmospheres and surface environments, but its actual contribution is not well justified. Young Sun-like (G/K-type) stars, particularly those younger than 600 Myr, exhibit extremely active magnetic behaviour, producing “superflares" with energies exceeding erg nearly daily, as revealed by Kepler and TESS observations (e.g. Maehara et al., 2012; Namekata et al., 2021). This suggests that such energetic flares on the early Sun could have significantly impacted atmospheric escape and chemistry on early Venus, Earth, and Mars when life may have first emerged. Likewise, exoplanets in the habitable zones of M dwarfs orbit very close to their host stars and may be subjected to repeated flare-driven UV irradiation, potentially altering their atmospheric chemistry (Chen et al., 2021). Thus, understanding the UV–optical continuum radiation of flares across stellar types is central to both exoplanetary habitability and early solar system evolution.
@@
Objective 1: AU Mic presents a unique opportunity to characterise the continuum properties of M-dwarf flares. Thanks to Mauve’s seamless 200–700 nm wavelength coverage, a time-resolved UV–optical spectra can be obtained. The cumulative flare frequency distribution of AU Mic derived from TESS Cycles 1 and 3 (Figure 2; Ikuta et al., 2023) suggests that a 10-day observation should detect at least one flare with a TESS-band amplitude of 2.4% and several events with amplitudes around 1%. Since flares with amplitudes 0.1% occur at a rate of 3 day-1 and typically last less than 30 minutes, over 90% of time-resolved spectra will be obtained during quiescent periods. By summing more than 360 quiescent spectra (corresponding to 1800 sec exposure), a high S/N pre-flare spectrum will be constructed to enable reliable flare-only spectral extraction. Simulated Mauve spectra of a flare with a 300-sec exposure indicate that, after binning over 40–80 pixels (10–20 Å), the blackbody continua with and 20,000 K, as well as continua composed of a 10,000 K blackbody plus an optically thin Balmer-jump component with a Balmer jump ratio of 3, can be distinguished. See Figure 3 for the model calculation. The minimum success criterion of this theme is to detect at least one flare and measure its effective temperature and Balmer jump ratio, while full success consists of detecting 3–4 flares and investigating the diversity of their NUV continua. If a large superflare occurs by chance, Mauve’s short exposure capability (30–60 sec) will allow us to resolve temporal changes in the continuum throughout the flare, providing unique constraints on the heating processes governing powerful M-dwarf flares.
@@
Objective 2: Young Sun-like (G/K-type) stars such as V889 Her and Eridani offer an equally important window into the continuum properties of stellar superflares. These stars are rapid rotators with strong magnetic fields and produce frequent superflares with energies erg, but the associated brightening in the optical band is only 0.5–1%, making ground-based spectroscopic observations nearly impossible. TESS observations indicate that V889 Her flares of this scale occur every 1–2 days and last tens of minutes (Namekata et al., 2025). Simulations assuming a 1% flare ( erg) and a 5-minute Mauve exposure show that spectra binned to 25–50 nm can distinguish optically thin continua—characterised by a pronounced Balmer jump—from 10,000–20,000 K blackbody continua. Thus, observations are planned for V889 Her for 100–380 hrs considering the high flare rate in TESS. Eridani shows only a modest flare rate in TESS photometry, yet HST FUV observations reveal that it produces flares far more frequently (Loyd et al., 2022). Although Mauve’s sensitivity is more limited than that of HST, the largest flares detectable by Mauve are expected to occur roughly once or twice per day. Therefore, the plan is to conduct 50–100 hours of observations of Eridani to capture these events.
@@ 4.3.2 Probing coronal mass ejections through UV dimming signatures
A particularly valuable solar diagnostic is the He II 30.4 nm line, formed in the transition region at K. Sun-as-a-star analyses with SDO/EVE and GOES/EUVS have shown that this line exhibits measurable obscuration dimmings during filament eruptions leading to coronal mass ejections (Xu et al., 2024). This dimming corresponds to the transit of erupting filament material across the solar disk, which significantly attenuates the background emission. The amplitude and duration of this event demonstrate the sensitivity of transition-region diagnostics to CME-related obscuration and provide a benchmark for what could be detectable in stellar light curves. Beyond the transition region, dimming has also been observed in chromospheric diagnostics: for instance, Jiang et al. (2007) reported H dimmings associated with filament eruptions, further supporting the notion that multi-wavelength dimming phenomena are robust CME tracers across atmospheric layers. Together, these results reinforce the physical link between post-flare dimming signatures and CME mass budgets.
Building on this solar benchmark, Mauve’s high-cadence spectrophotometry offers a unique opportunity to extend dimming-based CME detection to other stars. While the He II 30.4 nm line cannot be detected by Mauve, it is a suitable proxy for dimming tracers within stellar chromospheres. Several spectral diagnostics within Mauve’s bandpass probe chromospheric and transition-region layers at – K, making them conceptually analogous to He II 30.4 nm. These include the Mg II h/k resonance lines (2796/2803 Å), the Ca II H&K lines (3934/3968 Å), and the Balmer series (e.g., H 6563 Å, H 4861 Å). Each is sensitive to plasma density variations and obscuration by cool erupting material, with line cores expected to show the strongest dimming signatures. By monitoring these diagnostics after energetic flares, Mauve can search for persistent flux decreases beyond the flare’s impulsive phase. The wavelength dependence and timing will make such signatures compelling indirect evidence of stellar coronal mass ejections.
@@ 4.3.3 Quiescent UV Emission in Low-Mass Stars
To constrain the evolution of rotation and magnetic activity in Sun-like (and lower-mass) stars, a sample of stars in open clusters with well-determined ages will be used (Agüeros_2011; Agüeros et al., 2018; Curtis et al., 2019b, a, 2020; Douglas et al., 2014; Núñez and Agüeros, 2016; Núñez et al., 2017, 2015; Núñez_2022; Núñez et al., 2024).
@@
Mauve data of proposed targets will allow the exploration into how UV emission varies across late F and G stars with different activity levels, rotation rates, and ages. Furthermore, several of these targets also appear in the Habitable Worlds Observatory Preliminary Input Catalog. Although the focus is on the connection between rotation, activity, and age, these targets underscore the broader value of this work. Improving the understanding of the high-energy environments of solar-type stars is essential for interpreting future Habitable Worlds Observatory observations of planetary systems and their atmospheres.
@@ 4.4.1 Young Exoplanet Hosts and Habitable Worlds Observatory
By providing visible to NUV spectrophotometry of these targets, Mauve can fill in gaps in the spectral energy distributions of these stars to inform Habitable Worlds Observatory on mission design requirements, yield estimates, and contribute to a library of panchromatic spectrophotometry for main-sequence stars. The minimum requirement is to obtain a single snapshot of the near-ultraviolet spectrum that can be published as a resource for any scientific investigation that requires the near-ultraviolet spectrum of a main sequence star as an input, but the primary intended purpose is to provide the Habitable Worlds Observatory Target Stars and Systems Working Group to inform the survey plan. One plausible scientific use for this library is to compare the obtained spectra to the various stellar atmosphere models in circulation, such as the PHOENIX models, to determine how they differ from ground truth data for different spectral types. Another is to inform planet atmosphere modelling codes that need near-ultraviolet spectra as inputs for work on haze production and biosignatures.

### v3

@@ Abstract
Pioneering a new model for science satellites and space science (Archer et al., 2020), Mauve launched on 28th November 2025 and will begin a collaborative scientific programme in early 2026, that focuses on monitoring stellar activity and variability. A major strength of the Mauve science programme is the availability of thousands of observational hours each year for its consortium members. This makes it ideal for time-domain astronomy, where hundreds of hours can be dedicated to both, long-duration observations and repeat observations. Additionally, Mauve will be ideally placed to conduct pilot studies and high-risk investigations.
@@ 2 Platform and Payload
As final operational performance will only be known post-commissioning, the BSSL team and the Mauve Science Team are considering multiple performance scenarios to prepare the science programme and optimise the observation schedule. MauveSim (Saba et al., 2025a), an end-to-end simulator, was developed to assess the science capabilities of the satellite. The simulations presented here are derived from the most recent ground-based testing and expected in-orbit performance, with appropriate margins applied. The detector temperature, for instance, directly influences the thermally induced noise on the detector and the system’s overall performance. Following commissioning, MauveSim will be revised and updated to reflect the final in-flight performance. Pending commissioning and final performance, some Mauve science themes may be updated to account for spacecraft in-orbit behaviour.
@@ 3 Commissioning and Calibration
Mauve is equipped with a low-resolution spectrometer, resulting in each detector resolution element spanning approximately 40 pixels. This configuration leads to significant oversampling of the instrument line function – a matrix that describes photon energy redistribution, i.e., the probability of detecting an incident photon at any given pixel. Due to this oversampling, each pixel collects photons from a range of wavelengths, making it impossible to determine the exact energy of a photon striking a specific pixel. Any observation obtained with Mauve can be represented as the matrix product of the true stellar spectrum, the telescope’s effective area (ARF, representing total detector throughput), and the instrument line function (RMF). However, this matrix operation is non-reversible, meaning that the true stellar spectrum cannot be directly separated from the instrumental effects to recover the intrinsic stellar flux in physical units. Because all flux calibration information is contained within the ARF and RMF, calibrating Mauve involves adjusting these arrays such that the observed spectrum of a stable reference star with a well-known spectrum matches the simulated spectrum of the same source. When the observed and simulated spectra of the reference star match with each other, the calibration is deemed successful. If discrepancies remain, the ARF and RMF are iteratively refined until a satisfactory match is achieved. During the commissioning phase and routinely throughout nominal operations, BSSL will generate and update the calibration files required by users to perform accurate data analysis.
The pointing stability was simulated and modelled by ISISpace, but is a primary analysis of the commissioning stage. Prior to launch the relative pointing error (RPE) of the system was estimated to be approximately 12” across a 10 s window. This information has been incorporated into a pointing model for the system, enabling time-domain jitter analysis. Our estimates indicate that jitter-induced noise is negligible. The fibre diameter is 230 m, corresponding to a half-cone FoV of approximately 47”.
@@
To establish the initial post-launch pointing calibration, we selected Jupiter due to its exceptional brightness and large angular size. The planet enabled a first-order determination of the initial misalignment between the star tracker attitude solution and the telescope boresight. Subsequently, bright point sources (stars) were used to confirm and refine the misalignment identified from the Jupiter observations.
A comprehensive commissioning and calibration report will be released following Mauve’s commissioning period, summarising the lessons learned during spacecraft commissioning and outlining the achievable calibration performance of the instrument. The commissioning phase is currently expected to last up to four months. However, BSSL will review this timeline with its industry partners and announce any schedule updates as needed.
@@ 4.1 Mauve Science Planning
While Blue Skies Space Ltd. is responsible for the development and operation of the satellite platform, as well as implementation of the observation schedule, the Mauve Survey Programme is designed as a collaborative, member-driven scientific initiative. Survey members contribute to the definition of science themes, objectives, target lists, and observational strategies through a structured planning process conducted in coordination with Blue Skies Space Ltd..
The programme is led by the Mauve Science Team, a consortium of scientists from institutions across Europe, Asia, and North America. The Science Team was established during the survey design phase, prior to Mauve launch, through a membership subscription model under which participating organisations joined the collaborative survey programme. Institutions that joined at this early stage also joined the Science Team and contributed to survey planning and prioritisation activities.
The participating member institutions include:
@@ 4.2 Science Overview
Pre-flight constraints on the field of regard (FoR) limit Mauve observations to targets within the declination range of +31.8° to -46.4°; sources outside these bounds are currently inaccessible. In addition, pre-flight analyses were performed to estimate the telescope sensitivity limits and the corresponding target magnitude thresholds across different wavelength bands. The sensitivity plots in Fig. 1 show the expected signal-to-noise ratio (S/N) per 10.5 nm bin (one resolution element) as a function of visible magnitude for several wavelength bands. These predictions are based on pre-launch performance models and ground-based characterisation of the dark current noise (450 counts s-1) and overall payload efficiency (see Saba et al. (2025b)). These sensitivity estimates are being progressively refined through ongoing commissioning activities and in-flight performance characterisation. Updated sensitivity curves, reflecting the measured on-orbit performance of the instrument, will be released following completion of the commissioning and calibration phase.
@@
The selected science themes can be broadly categorised into the following research areas: stellar activity and variability, host-star/planet interaction, hot stars, and exotic populations in binaries. Specific themes can then also be grouped by three main observation strategies, i.e. long-duration monitoring programmes, short-cadence single snapshots, and short-cadence monitoring with repeats. These categories determine scheduling strategy and observation cadence. Long-duration monitoring (10-250 hours) is typically focused on tracking stellar variability on specific targets over time. The themes using this observation strategy include studies of stellar activity and variability, specifically to understand the flare occurrence rates and to constrain the signatures of coronal mass ejections through flare-associated dimming.
The second category comprises short-cadence (<5 hours) single snapshots, typically associated with large-scale statistical programmes observing many targets, or small groups of targets observed repeatedly. The themes related to short-cadence single-snapshot observations include investigations of exoplanet host stars and future Habitable Worlds Observatory targets, the study of quiescent near-ultraviolet emission in low-mass stars to constrain the evolution of rotation and magnetic activity in Sun-like (and lower-mass) stars, and the classical Be star survey, designed to explore the mechanism that causes disk ejection.
@@ 4.3.1 Radiation mechanisms of stellar flares on M-dwarfs and young Sun-like stars
The UV and X-ray radiation from stellar flares is thought to influence planetary atmospheres and surface environments, but its actual contribution is not well justified. Young Sun-like (G/K-type) stars, particularly those younger than 600 Myr, exhibit extremely active magnetic behaviour, producing “superflares" with energies exceeding erg above their quiescent emission nearly daily, as revealed by Kepler and TESS observations (e.g. Maehara et al., 2012; Namekata et al., 2021). This suggests that such energetic flares on the early Sun could have significantly impacted atmospheric escape and chemistry on early Venus, Earth, and Mars when life may have first emerged. Likewise, exoplanets in the habitable zones of M dwarfs orbit very close to their host stars and may be subjected to repeated flare-driven UV irradiation, potentially altering their atmospheric chemistry (Chen et al., 2021). Thus, understanding the UV–optical continuum radiation of flares across stellar types is central to both exoplanetary habitability and early solar system evolution.
@@
Objective 1: AU Mic presents a unique opportunity to characterise the continuum properties of M-dwarf flares. Thanks to Mauve’s seamless 200–700 nm wavelength coverage, a time-resolved UV–optical spectra can be obtained. The cumulative flare frequency distribution of AU Mic derived from TESS Cycles 1 and 3 (Figure 3; Ikuta et al., 2023) suggests that a 10-day observation should detect at least one flare with a TESS-band amplitude of 2.4% and several events with amplitudes around 1%. Since flares with amplitudes 0.1% occur at a rate of 3 day-1 and typically last less than 30 minutes, over 90% of time-resolved spectra will be obtained during quiescent periods. By summing more than 360 quiescent spectra (corresponding to 1800 sec exposure), a high S/N pre-flare spectrum will be constructed to enable reliable flare-only spectral extraction.
Simulated Mauve spectra of a flare with a 300-sec exposure indicate that, after binning over 40–80 pixels (10–20 Å), the blackbody continua with Teff=10,000 and 20,000 K can be distinguished. In addition, the difference in the simulated spectra for the blackbody continua Teff=10,000 and for the continua composed of a 10,000 K blackbody plus an optically thin Balmer-jump component with a Balmer jump ratio of 3 suggests that the NUV excess flux in the wavelength shorter than the Balmer limit (346.6 nm) due to the Balmer-jump component can be detectable with Mauve. See Figure 4 for the model calculation. The minimum success criterion of this theme is to detect at least one flare and measure its effective temperature and Balmer jump ratio, while full success consists of detecting 3–4 flares and investigating the diversity of their NUV continua. If a superflare occurs by chance, Mauve’s short exposure capability (30–60 sec) will allow us to resolve temporal changes in the continuum throughout the flare, providing unique constraints on the heating processes governing powerful M-dwarf flares.
@@
Objective 2: Young Sun-like (G/K-type) stars such as V889 Her and Eridani offer an equally important window into the continuum properties of stellar superflares. These stars are rapid rotators with strong magnetic fields and produce frequent superflares with energies erg. Although the associated brightening in the optical band is only 0.5–1%, this does not reflect the total energetic output of the event. While cool stars have photospheric temperatures of roughly 3000-6000 K, the flare emission can be characterised by an effective temperature of 10,000 K (or higher). This naturally leads to a stronger relative enhancement at shorter wavelengths (UV and X-rays), while the optical contributes only a small fraction to the bolometric flare energy. The low optical contrast makes ground-based spectroscopic observations nearly impossible. TESS observations indicate that V889 Her flares of this scale occur every 1–2 days and last tens of minutes (Namekata et al., 2025). Simulations assuming a 1% increase in flux relative to the quiescent emission ( erg) and a 5-minute Mauve exposure show that spectra binned to 25–50 nm can distinguish optically thin continua—characterised by a pronounced Balmer jump—from 10,000–20,000 K blackbody continua. Thus, observations are planned for V889 Her for 100–380 hrs considering the high flare rate in TESS. Eridani shows only a modest flare rate in TESS photometry, yet HST FUV observations reveal that it produces flares far more frequently (Loyd et al., 2022). Although Mauve’s sensitivity is more limited than that of HST, the largest flares detectable by Mauve are expected to occur roughly once or twice per day. Therefore, the plan is to conduct 50–100 hours of observations of Eridani to capture these events.
@@ 4.3.2 Probing coronal mass ejections through UV dimming signatures
A particularly valuable solar diagnostic is the He II 30.4 nm line, formed in the transition region at K. Sun-as-a-star analyses with SDO/EVE and GOES/EUVS have shown that this line exhibits measurable obscuration dimmings during filament eruptions leading to coronal mass ejections (Xu et al., 2024). Figure 5 shows that, following an eruptive flare, the Sun-as-a-star (full-disk integrated) light curve of the He II 30.4 nm line exhibits a CME-associated dimming. This dimming corresponds to the transit of erupting filament material across the solar disk, which significantly attenuates the background emission. The amplitude and duration of this event demonstrate the sensitivity of transition-region diagnostics to CME-related obscuration and provide a benchmark for what could be detectable in stellar light curves. Beyond the transition region, dimming has also been observed in chromospheric diagnostics: for instance, Jiang et al. (2007) reported H dimmings associated with filament eruptions, further supporting the notion that multi-wavelength dimming phenomena are robust CME tracers across atmospheric layers. Together, these results reinforce the physical link between post-flare dimming signatures and CME mass budgets.
Building on this solar benchmark, Mauve’s high-cadence spectrophotometry offers a unique opportunity to extend dimming-based CME detection to other stars. While the He II 30.4 nm line cannot be detected by Mauve, it is a suitable proxy for dimming tracers within stellar chromospheres. Several spectral diagnostics within Mauve’s bandpass probe chromospheric and transition-region layers at – K, making them conceptually analogous to He II 30.4 nm (e.g. (Veronig et al., 2025; Namekata et al., 2021; Doyle and Collier Cameron, 1990)). These include the Mg II h/k resonance lines (2796/2803 Å), the Ca II H&K lines (3934/3968 Å), and the Balmer series (e.g., H 6563 Å, H 4861 Å). Each is sensitive to plasma density variations and obscuration by cool erupting material, with line cores expected to show the strongest dimming signatures. By monitoring these diagnostics after energetic flares, Mauve can search for persistent flux decreases beyond the flare’s impulsive phase. The wavelength dependence and timing will make such signatures compelling indirect evidence of stellar coronal mass ejections.
@@ 4.3.3 Quiescent UV Emission in Low-Mass Stars
To constrain the evolution of rotation and magnetic activity in Sun-like (and lower-mass) stars, a sample of stars in open clusters with well-determined ages will be used (Agüeros_2011; Agüeros et al., 2018; Curtis et al., 2019b, a, 2020; Douglas et al., 2014; Núñez and Agüeros, 2016; Núñez et al., 2017, 2015; Núñez_2022; Núñez et al., 2024). We will use Mauve to capture short-duration (i.e., minutes or hours) variability, typically associated with flares. Such flares will then be removed from our UV emission measurement to capture the quiescent level of emission.
@@
Mauve data of proposed targets will allow the exploration into how UV emission varies across late F and G stars with different activity levels, rotation rates, and ages. Furthermore, several of these targets also appear in the Habitable Worlds Observatory Preliminary Input Catalog (Mamajek and Stapelfeldt, 2024). Although the focus is on the connection between rotation, activity, and age, these targets underscore the broader value of this work. Improving the understanding of the high-energy environments of solar-type stars is essential for interpreting future Habitable Worlds Observatory observations of planetary systems and their atmospheres.
@@ 4.4.1 Young Exoplanet Hosts and Habitable Worlds Observatory
By providing visible to NUV spectrophotometry of these targets, Mauve can fill in gaps in the spectral energy distributions of these stars to inform Habitable Worlds Observatory on mission design requirements, yield estimates, and contribute to a library of panchromatic spectrophotometry for main-sequence stars. The minimum requirement is to obtain a single snapshot of the near-ultraviolet spectrum that can be published as a resource for any scientific investigation that requires the near-ultraviolet spectrum of a main sequence star as an input, but the primary intended purpose is to provide the Habitable Worlds Observatory Target Stars and Systems Working Group to inform the survey plan. One plausible scientific use for this library is to compare the obtained spectra to the various stellar atmosphere models in circulation, such as the PHOENIX models, to determine how they differ from ground truth data for different spectral types. Another is to inform planet atmosphere modelling codes that need near-ultraviolet spectra as inputs for work on haze production, chemical disequilibrium, and biosignatures.